# Adversarial Evasion Attacks Practicality in Networks: Testing the Impact of Dynamic Learning Authors’ draft for soliciting feedback: April 4, 2024

### v3

@@ II-C2 Defenses Against Adversarial Attacks
III Related Work
III-A Adversarial attacks against ML-NIDS
Numerous researchers have examined adversarial attacks against machine learning-based network intrusion detection systems. Alatwi et al. [23] surveyed and categorized the studies of adversarial attacks against ML-based NIDSs and briefly discussed their applicability in real-world scenarios. He et al. [24] has also conducted a comprehensive exploration of white- and black-box adversarial attacks on DNNs, and their potential relevance in NIDS. They also analyzed existing defense mechanisms against these attacks. Furthermore, Martins et al. [25] examined adversarial machine learning against intrusion and malware detection systems. They concluded that a wide variety of the reviewed attacks could be effective, however, their practicality was not been tested enough in intrusion scenarios. Furthermore, They have stated that some of the datasets in intrusion scenarios are outdated, and most of them are similar and lack variety.
III-B Practicality of Adversarial attacks against ML-NIDS
Some studies have focused on the practicality of adversarial attacks in the network intrusion detection domain; Vitorino et al. [26] presented an overview of adversarial learning techniques capable of generating realistic examples to deceive NIDSs. Apruzzese et al. [4] aimed to specify and model the abilities and circumstances necessary for attackers to execute practical adversarial attacks on network intrusion detection systems. Both [26] and [4] primarily reviewed and evaluated attacks from research literature but did not consider the dynamic nature of modern ML models. In a separate study, Apruzzese et al. [3] emphasized the existence of a gap between researchers and practitioners in the field of adversarial attacks. Their primary focus was not on NIDSs; however, through collaboration between industry and academia, they articulated the general viewpoint that "Real attackers don’t compute gradients." In other words, they were unable to identify an attacker who computes gradients to craft real-world adversarial attacks.
Merzouk et al. [5] attempted to inspect the practicality of adversarial attacks to evade network intrusion detection systems. They highlighted some limitations of adversarial attacks against ML-NIDS. However, they have only concentrated on the feature space attacks.
To the best of our knowledge, none of the previous work has explored and tested the practicality of adversarial attacks against dynamic ML models in the intrusion detection domain.
@@ III-B Practicality of Adversarial attacks against ML-NIDS
for tree= draw, text centered [ Adversarial
Evasion Attack
Against ML-NIDS [Create Adversarial
Packets that produces
perturbed features
(Problem-Space), angle below [Access Network
Packets, name =y] [Add perturbations
on Network Packets, angle below [Maintain Network
Functionality] [Maintain Malicious
Functionality
?] [Reverse Engineer
the packets in
a way that
produces perturbed
features after
the features
extraction, angle below [Pick perturbable
Features[Predict some
Feature Vector Items] [Insider
Knowledge
(White-Box)
?]][Know
up-to-date
information of
the model, name =x [Insider
Knowledge
(White-Box)
?] [Query
the model
(Black-Box) [Access Model’s
input and
output
?]]]]]] [Adversarial
Features
(Feature-Space), angle below [Access Feature
Vector
?] [Add perturbations
on Perturable
Features, angle below[Pick perturbable
Features[Predict some
Feature Vector
Items] [Insider
Knowledge
(White-Box)
?]] [Know
up-to-date
information of
the model [Insider
Knowledge
(White-box)
?] [Query
the model
(Black-Box) [Access Model’s
input and
output
@@
IV Identifying Infeasible Adversarial Attack Prerequisites Through an Attack Tree Threat Model
The majority of the adversarial attacks were designed and tested in the literature using computer vision datasets [3]. Because of the focus on the computer vision datasets, it is not clear if these attacks will be applicable in the same way to the network/cybersecurity domain which has more adversaries and more security constraints and measures, making it much more complicated to evade. This contrast is depicted in Figure 2, which illustrates the crafting of evasion adversarial attacks in both the computer vision and network domains. Thus, in this section, we present an attack tree, depicted in Figure 3, that showcases the various strategies attackers might employ to execute evasion adversarial attacks. Additionally, we identify leaf nodes (prerequisites) with questionable feasibility marked by the ’?’ sign.
The attack tree presented in Figure 3 outlines an Adversarial Evasion Attack strategy targeting Machine Learning-based Network Intrusion Detection Systems (ML-NIDS).
The tree diverges into two main branches, representing the two alternative ways attackers can fool the model: Adversarial Features (Feature space) and Adversarial Packets (Problem-Space). The right branch focuses on the manipulation of feature vectors to fool ML-NIDS. To achieve this, attackers require access to the feature vector and the ability to introduce perturbations to specific features. However, directly accessing the feature vector presents a significant obstacle for external attackers, rendering this approach highly improbable. On the other hand, introducing perturbations to the feature vector necessitates identifying perturbable features (not all features are susceptible to manipulation, as will be discussed further) and requires up-to-date knowledge of the underlying ML-NIDS model. Identifying perturbable features necessitates information about the specific features employed by the target ML-NIDS. Acquiring this information through insider knowledge is highly unlikely. An alternative option for the attackers is to make informed assumptions or predictions regarding certain features, an approach subject to uncertainties. As for up-to-date knowledge of the model, the attacker could get it by insider knowledge (questionable practicality) or by querying the model. However, as illustrated in Figure 1, accessing the model’s input and output is nearly impossible for an outsider. This presents a significant challenge for attackers attempting to launch adversarial attacks against ML-NIDS.
In the left branch of the tree, we delve into Problem-Space attacks, which involve manipulating network packets to evade ML-NIDS detection. In this scenario, the attacker needs to gain access to network packets and introduce perturbations to them. Accessing network packets is comparatively more feasible than accessing the feature vector of the model, making problem-space attacks appear more practical than feature-space attacks. However, adding perturbations to network packets requires the attacker to preserve network functionality, maintain malicious functionality (since it’s an attack), and reverse engineer the packets to produce perturbed features after extraction. We find maintaining the malicious functionality of network packets presents an uncertain feasibility, as existing adversarial attacks against ML-NIDS lack a comprehensive evaluation of the maliciousness of their generated examples. These attacks primarily focus on evading detection by the model, neglecting the preservation of the intended malicious behavior [24]. Additionally, the process of reverse engineering the packets entails similar requirements as the feature-space attack outlined in the right branch of the tree.
The attack tree illustrates that achieving the root node goal is highly challenging. Although problem-space attacks may seem more practical, we find that both problem and feature-space attacks have limited suitability in real-world scenarios.
@@ IV Identifying Infeasible Adversarial Attack Prerequisites Through an Attack Tree Threat Model
for tree= draw, text centered [ Practicality Issues of Adversarial
Evasion Attacks Against ML-NIDS [Attackers’
Knowledge[White-Box [Own
ML-NIDS [Very unlikely
for an
attacker to
get knowledge
of the model]] [Bought
Pre-built
ML-NIDS [Different settings
at deployment
still Very
unlikely
for an attacker
to get knowledge
of the model]]] [Black-Box [Basic
Knowledge][Querying
NIDSs]]] [Attack’s
Space [ Feature
Space
Perturbations [Accessing the
Feature Vector] [NIDS
Features ,name=x [Imperturbable
Features] [Correlated
Features] [Heterogeneous
Features] [Few remaining
Perturbable
Features]]] [ Problem
Space
Perturbations, name =y [Network
Functionality][Malicious
Functionality][Inverse
Feature-
Mapping
Problem] [Side-
Effect
Features]]] [Dynamic
ML [Attacker
must
always have
up-to-date
knowledge
of the model
which is
[…]
@@
V Taxonomy of Practicality Isues of Adversarial Attacks Against ML-NIDS
V-A Attackers’ Knowledge
V-A1 White-Box Feasibility
V-A2 Black-Box Feasibility
As mentioned in section II, attackers in black-box settings have no knowledge of the target ML model. There is no doubt that black-box evasions are the more practical type of adversarial attack [23] as obtaining all the needed information about the target model is extremely difficult in the network security domain. In the next sub-section, we will explore the feasibility of black-box attacks against ML-based NIDS.
Basic Knowledge
In recent years, researchers successfully crafted adversarial examples without previous knowledge of the target’s learning algorithm, gradients, parameters, training data, or hyper-parameters [27]. However, only a minority of these researches use intrusion, malware, or spam detection datasets. As seen in figure 2a, to generate adversarial examples in the computer vision domain, the attackers adds perturbation by slightly changing the pixels’ values. The medium in the computer vision domain is well-known as the attackers are perturbing a digital image that is visible to them. On the other hand, in network intrusion, the attackers don’t know the medium they are trying to perturb. Each ML-based network intrusion detection system uses its unique pre-processing/feature extraction and selection techniques. If the attackers don’t know the selected features and the pre-processing methods of the target model, then they don’t know what to perturb. Kuppa et al. [28] labeled their attack as black-box, however, they assumed that the adversaries have previous knowledge of the target model’s feature set, feature extractor, and feature transformers. We found that most of the "black-box" attacks in the NIDS domain make similar assumptions. In our opinion, if the attackers know all this information about the features management and selection of the target model, their attack cannot be considered black-box.
Querying NIDS
As mentioned in section II, one of the main corners of black-box adversarial attacks is querying the target model to obtain a substitute model. Then using that alternative model, the attackers can get all the knowledge they need to craft their adversarial attack. It would be reasonable to assume that only the system management teams have access to the output of NIDS, as seen in Figure 1. Thus, it is not clear how querying NIDSs in black-box attack settings is possible if the attacker is not an insider.
V-B Attack’s Space
V-B1 Feature-Space Perturbations
@@
Accessing the Feature Vector:
Normally, attackers don’t have access to the feature vector after the feature extraction phase in the ML-based NIDS pipeline. So how would feature-space adversarial attacks against NIDS be considered practical? That is an important question to answer considering that most of NIDS’s adversarial attack research is in feature-space attacks [4].
NIDS Features:
As previously mentioned, knowing which features, and feature transformation techniques are used by the target model is a very hard task for the attacker. However, even if the attackers know the selected features and have access to them, adding perturbations to the ML-based NIDS features is a complicated task [5]. Unlike the computer vision domain, numerous constraints make crafting adversarial attacks in the NIDS space tough.
  • Imperturbable Features: A sizable portion of features in the NIDS datasets are boolean (such as flags) and discrete numbers (such as the protocol feature). The presence of these features doesn’t give flexibility to the attackers to add perturbations that maximizes the loss to fool the target model. For example, the attacker cannot add a perturbation to a flag in the positive direction if its value is 1 (because the maximum value is 1). The same goes for a discrete feature, for instance, the attacker cannot add an arbitrary value to the protocol feature because some protocol numbers are unassigned (e.g., unassigned IP address). Moreover, some numerical values have similar limitations, for example, the attacker cannot change the packet length feature to a negative value.
  • Correlated Features: NIDS typically extract various features from network traffic to identify malicious activity. These features are not always independent and can exhibit dependencies or correlations between them. For example, TCP flags in packets and TCP as the transport protocol are correlated [29]. Unlike in computer vision, in the NIDS domain, attackers cannot simply add perturbations to independent pixels; instead, they must consider the complex relationships between different features, which complicates the crafting of attacks in the NIDS domain.
@@
  • Few Remaining Perturbable Features: As mentioned in Section II, feature reduction can be a potential defense strategy against adversarial attacks, as it reduces the attack surface and increases the model’s robustness. Due to the NIDS feature constraints mentioned in this section, the NIDS domain has fewer perturbable features compared to other domains like computer vision, which act as a natural line of defense against adversarial attacks.
V-C Problem-Space Perturbations

### v5

@@ 4.3.2 Defenses Against Adversarial Attacks
The attack tree presented in Figure 7 outlines an Adversarial Evasion Attack strategy targeting Machine Learning-based Network Intrusion Detection Systems (ML-NIDS). The tree diverges into two main branches, representing the two alternative ways attackers can fool the model: Adversarial Features (Feature space) and Adversarial Packets (Problem-Space). The right branch focuses on the manipulation of feature vectors to deceive ML-NIDS. To achieve this, attackers require access to the feature vector and the ability to introduce perturbations to specific features. However, directly accessing the feature vector presents a significant obstacle for external attackers, making this approach highly impractical. Introducing perturbations to the feature vector requires identifying perturbable features (not all features are susceptible to manipulation, as will be discussed later) and requires up-to-date knowledge of the underlying ML-NIDS model. Identifying perturbable features requires information about the specific features employed by the target ML-NIDS. Acquiring this information through insider knowledge is highly unlikely. An alternative option for attackers is to make informed assumptions or predictions regarding certain features, an approach fraught with uncertainties. As for up-to-date knowledge of the model, the attacker could obtain it through insider knowledge (questionable practicality) or by querying the model directly. However, as illustrated in Figure 2, accessing the model’s input and output is very improbable for an outsider, as NIDSs are not designed to provide feedback when queried [38]. Nonetheless, side-channel querying represents another potential avenue. In essence, the fundamental limitation of directly accessing the feature vector renders the right branch (feature-space attacks) highly impractical for real-world scenarios.
In the left branch of the tree, we delve into problem-space attacks, which involve manipulating network packets to evade ML-NIDS detection. In this scenario, the attacker needs to gain access to network packets and introduce perturbations to them. Accessing network packets is comparatively more feasible than accessing the feature vector of the model, making problem-space attacks appear more practical than feature-space attacks. However, adding perturbations to network packets requires the attacker to preserve network functionality, maintain malicious functionality (since it’s an attack), handle side-effect features (a.k.a. collateral damage features) [24] [27], and reverse engineer the packets to produce perturbed features after extraction, in addition to addressing all the constraints presented in the right branch (feature-space attacks), since the main aim of problem-space attacks is to add perturbations to features after feature extraction. The tree shows that maintaining network functionality and malicious functionality are required simultaneously (AND condition), while handling side-effect features represents an additional constraint that attackers must address when modifying network packets. The reverse engineering process presents multiple pathways with their own constraints. We find that maintaining the malicious functionality of network packets presents uncertain feasibility, as existing adversarial attacks against ML-NIDS lack a comprehensive evaluation of the maliciousness of their generated examples. These attacks primarily focus on evading detection by the model, neglecting the preservation of the intended malicious behavior [9]. Additionally, the process of reverse engineering the packets entails similar requirements as the feature-space attack outlined in the right branch of the tree, including the challenging aspects of obtaining model knowledge and identifying perturbable features.
The attack tree illustrates that evasion adversarial attacks against ML-NIDS remain highly challenging. Although problem-space attacks may appear more practical than feature-space approaches, both attack vectors face significant real-world constraints. However, the red branches in our attack tree indicate certain aspects that might possess some feasibility, which we will explore in subsequent sections to provide a comprehensive assessment of the attack landscape.
@@
6 Taxonomy of Practicality Isues of Adversarial Attacks Against ML-NIDS
6.1 Attackers’ Knowledge
6.1.1 White-Box Feasibility
6.1.2 Black-Box Feasibility
As mentioned in section 4, attackers in black-box settings have no knowledge of the target ML model. There is no doubt that black-box evasions are the more practical type of adversarial attack [10] as obtaining all the needed information about the target model is extremely difficult in the network security domain. In the next sub-section, we will explore the feasibility of black-box attacks against ML-based NIDS.
  • Preliminary Knowledge: Even under black-box assumptions, where the attacker lacks knowledge of the model itself (e.g., architecture, parameters, gradients) and relies on transfer-based attacks or random perturbations, some preliminary knowledge remains necessary in the NIDS domain, particularly regarding pre-processing, feature extraction, and feature selection. As shown in Figure 6a, in computer vision, adversarial examples are generated by slightly modifying pixel values, and the medium is well understood because attackers perturb pixels in a digital image (or add physical perturbations to physical objects, such as a stop sign). By contrast, in the problem-space of network intrusion, the medium to be perturbed is often unknown to the attacker. ML-based NIDSs employ distinct pre-processing pipelines and feature engineering choices. Without awareness of the target model’s selected features and pre-processing methods, attackers do not know what to perturb or how to perturb it. As shown in Figure 4, problem-space perturbations must produce corresponding feature-space perturbations after pre-processing on the selected features; thus, knowledge of these aspects is essential. Moreover, many works present black-box attacks while assuming knowledge of the target’s feature set, feature extractor, or feature transformers (e.g., Kuppa et al. [39]). In our view, when attackers possess such detailed information about feature management and selection, the attack no longer qualifies as black-box.
    However, as stated in our attack tree, some attackers might attempt to “predict some feature vector items” by assuming that certain features are selected based on research and assessment of commonly used and popular NIDS domain datasets. While this approach might be practical in some scenarios, there is no guarantee that the target ML-NIDS actually uses those predicted features. Furthermore, ElShehaby et al. [27] introduced a defense mechanism that specifically counters this approach by selecting only features that are not perturbable or hard to perturb in the problem space using their Perturb-ability Score (PS), which might render the assumption of “Predict some feature vector items” inapplicable.
  • Querying NIDS: As mentioned in section 4, one of the main corners of black-box adversarial attacks is querying the target model to obtain a substitute model. Then using that alternative model, the attackers can get all the knowledge they need to craft their adversarial attack. It would be reasonable to assume that only the system management teams have access to the output of NIDS, as seen in Figure 2. Thus, it is not clear how querying NIDSs in black-box attack settings is possible if the attacker is not an insider. Debicha et al. clearly stated that NIDSs are not designed to provide feedback [38].
    However, some threat models in adversarial attack research assume a powerful attacker might gain access to the NIDS system and control it to obtain input-output pairs for querying. This assumption raises a critical question: if an attacker possesses such extensive system access and control, why would they need to craft complex gradient-based evasion attacks when they could simply employ straightforward trial-and-error approaches? With direct access to the system, they could systematically test various perturbations in all directions, observe the system’s responses, and iteratively refine their attacks based on what proves effective, eliminating the need for sophisticated gradient computations or complex mathematical optimization techniques. We argue that if the attacker possesses that much power, then it is no longer an ML/adversarial attack problem anymore.
@@ 6.1.2 Black-Box Feasibility
6.2 Attack’s Space
6.2.1 Feature-Space Perturbations
@@
Accessing the Feature Vector: Normally, attackers don’t have access to the feature vector after the feature extraction phase in the ML-based NIDS pipeline. So how would feature-space adversarial attacks against NIDS be considered practical? That is an important question to answer considering that most of NIDS’s adversarial attack research is in feature-space attacks [4].
NIDS Features: As previously mentioned, knowing which features, and feature transformation techniques are used by the target model is a very hard task for the attacker. However, even if the attackers know the selected features and have access to them, adding perturbations to the ML-based NIDS features is a complicated task [5]. Unlike the computer vision domain, numerous constraints make crafting adversarial attacks in the NIDS space tough.
  • Imperturbable/Hard-to-Perturb Features: A sizable portion of features in the NIDS datasets are boolean (such as flags) and discrete numbers (such as the protocol feature). The presence of these features doesn’t give flexibility to the attackers to add perturbations that maximize the loss to fool the target model. For example, the attacker cannot add a perturbation to a flag in the positive direction if its value is 1 (because the maximum value is 1). The same goes for a discrete feature; for instance, the attacker cannot add an arbitrary value to the protocol feature because some protocol numbers are unassigned. Moreover, some numerical values have similar limitations; for example, the attacker cannot change the packet length feature to a negative value. Furthermore, perturbing some features in problem-space can totally change the functionality of the network flow, which can make the flow lose its malicious or network functionality, such as altering the destination IP [27]. Moreover, attackers cannot access some features such as backward features (from destination to source) [27], which represent the target system’s response characteristics including backward packet count, backward bytes, and backward inter-arrival times, as these are determined by the receiving system’s behavior rather than the attacker’s input.
  • Correlated Features: NIDS typically extracts various features from network traffic to identify malicious activity. These features are not always independent and usually exhibit dependencies or correlations among them. For example, TCP flags in packets and TCP as the transport protocol are correlated [40]. These correlations constrain attackers; for example, an attacker cannot simultaneously increase the flow duration while decreasing both the forward and backward inter-arrival times [27]. Unlike in computer vision, in the NIDS domain, attackers cannot simply add perturbations to independent pixels; instead, they must consider the complex relationships between different features, which complicates the crafting of attacks in the NIDS domain.
@@
  • Few Remaining Perturbable Features: As mentioned in Section 4, feature reduction can be a potential defense strategy against adversarial attacks, as it reduces the attack surface and increases the model’s robustness. Due to the NIDS feature constraints mentioned in this section, the NIDS domain has fewer perturbable features compared to other domains like computer vision, which act as a natural line of defense against adversarial attacks.
6.2.2 Problem-Space Perturbations
@@
6.3 Adaptivity to Dynamic ML
7 “Practical” Attacks
It is well established that a practical evasion adversarial attack must be black-box, problem-space, while maintaining malicious and network functionality. In addition to these basic prerequisites, based on our literature review, threat modeling (attack tree), and practicality issues taxonomy, we identified three potential nodes or ways to conduct a practical evasion adversarial attack on ML-NIDS: (1) Adding random or blind perturbations, (2) side-channel querying, and (3) knowing up-to-date information on training data of the target model to create a surrogate model This section explores attacks from literature that use these approaches to conduct evasion adversarial attacks on ML-NIDS. We also explore in this section whether they are just practical or whether there are some catches.
7.1 “Practical” Attacks from Literature
7.1.1 Adding Random or Blind Perturbations
Apruzzese et al. [42] present what we consider a practical evaison adversarial attack because it operates under realistic black-box constraints with no model knowledge, uses problem-space perturbations on raw packets, and maintains network/malicious functionality. The attack creates blind perturbations by adding 1-100 random bytes of padding to UDP packets and TCP packets with PSH flags using scapy, ensuring protocol compliance through proper checksum recreation. This approach is practical because it reflects the most constrained attacker scenario: infected hosts within an organization that attackers have compromised via exploits or accidental infection to generate malicious network traffic for objectives like exfiltration or sabotage. These attackers attempt evasion without sophisticated ML-NIDS knowledge while representing physically realizable attacks executable by any attacker controlling network communications.
Table 4: Summary of “Practical” Evasion Adversarial Attacks Against Network Intrusion Detection Systems
Paper & Authors Attack Type Year Black-box Level Problem-space NIDS Output Access Malicious Function Preserved Network Function Preserved Practical Constraints Adversarial Perturbations meet Concept Drift [42] Blind perturbations under concept drift 2024 COMPLETE - No model knowledge, unaware of drift YES - Raw packet perturbations NONE - Blind without feedback YES - Attack functionality maintained YES - Protocol compliance preserved Most restrictive: no model knowledge, no drift awareness, real-world deployment Vulnerability Disclosure through Adaptive Black-Box Attacks on NIDS [43] Adaptive with side-channel analysis 2025 COMPLETE - No architecture, parameters, or training data access YES - Metadata manipulation within protocol ranges SIDE-CHANNEL - CPU, memory, response times only YES - Attack packet content preserved YES - Protocol adherence maintained Metadata-only manipulation, Variance Inflation Factor (VIF) filtering preserves correlations, stealth operation Adv-Bot: Realistic Adversarial Botnet Attacks [38] Substitute model-based botnet attacks 2023 COMPLETE - Very limited target NIDS knowledge YES - Acts on traffic space respecting domain constraints SUBSTITUTE MODEL - No queries, uses surrogate models YES - All intended malicious functionality maintained YES - Domain constraints respected, valid traffic Surrogate models trained independently, transferability property exploitation, no target interaction
7.1.2 Side-Channel Querying
Ennaji et al. [43] present another practical route for evasion adversarial attacks. Their attack operates under strict black-box constraints using only passive side-channel observations like response times, CPU usage, and packet drop rates to infer NIDS behavior. The attack creates realistic perturbations through random walk modifications applied to network metadata (packet length, timestamps, ports) captured via network sniffing tools, ensuring protocol compliance. This approach is practical because it reflects genuine attacker capabilities through silent probing and gradual, natural traffic variations that blend with normal network activity, combined with adaptive feature selection using change-point detection and causality analysis to identify sensitive features without direct system access.
7.1.3 Knowing up-to-date information on training data of the target model to create a surrogate model
Debicha et al. [38] also introduced an evasion adversarial attack that operates under realistic black-box constraints with no knowledge of the target NIDS architecture or parameters, using only traffic-based manipulation that directly modifies raw network packets rather than extracted features. The attack achieves practicality through transferability-based evasion, where attackers train surrogate models using sniffed network traffic from their infected botnet machines, then generate adversarial botnet traffic that transfers to the target NIDS without requiring queries or feedback. The approach is realistic because it respects domain constraints through syntactic and semantic validation, manipulating only three directly controllable network factors: duration, packet count, and byte quantities. This is accomplished via time manipulation, packet injection/retention, and byte expansion/compression techniques using tools like Scapy and Hping, while maintaining the underlying malicious functionality of botnet communications.
Table 4 shows a summary of the “Practical” Evasion Adversarial Attacks Against Network Intrusion Detection Systems mentioned in this section.
7.2 The Catch?
Although we find that these attacks are the most practical, we still think that there is a catch in the practicality of these attacks and some hurdles that attackers must overcome.
For blind perturbations, they can actually backfire, making adversarial NetFlows easier to detect than their unperturbed malicious variants. Since attackers have no model knowledge and no feedback from the ML-NIDS, there’s no guarantee that perturbations will help with evasion and may actually harm their objectives. The attack’s effectiveness varies significantly across different ML algorithms and architectures, with some perturbations showing no statistically significant impact on detection rates [27].
Regarding side-channel querying attacks, while practical, there are significant catches limiting effectiveness. The major limitation is dependency on unreliable side-channel indicators that vary based on network load, hardware configurations, and environmental factors beyond attacker control. Additionally, the attack requires extensive computational overhead for change-point detection and causal analysis, making it time-consuming. Moreover, as they require sustained network access over extended periods, this increases detection likelihood by network administrators or advanced monitoring systems.